OpenStudy (anonymous):

how many ounces of a 40% alcohol solution should be mixed with 54 oz of a 70 % alcohol solution to produce a 60 % alcohol solution ???

OpenStudy (anonymous):

let x ounces of 40% solution be added. \[then \frac{ 40x }{100 }+\frac{ 54\times70 }{ 100 }=\frac{ \left( 54+x \right)\times60 }{100 }\] \[40x+\left( 54\times70 \right)=\left( 54+x \right)\times60\] \[4x+\left( 54\times7 \right)=\left( 54+x \right)\times6\] \[4x+378=324+6x\] 6x-4x=378-324 2 x=54 x=54/2=27 ounces